OpenStudy (millsemily):

Can someone check my answers real quickly? A line passes through the point (4,-3) and has slope m=1. Find the equation in point-slope form. A. y - 4 = x - 3 B. y + 3 = x + 4 C. y + 4 = x - 3 D. y + 3 = x - 4 <-- My answer, is this correct? (Point-slope form is y-y1=m (x-x1)) I kinda think it might be, B as well but I'm unsure. Thanks :) Oh! If it's wrong can you tell me what I did wrong?

OpenStudy (texaschic101):

y - y1 = m(x - x1) (4,-3)...x1 = 4 and y1 = -3 slope(m) = 1 now we sub y - (-3) = m(x - 4) = y + 3 = m(x - 4) you are correct...it is D

OpenStudy (millsemily):

I should of just done that, I made it more complicated than it needed to be Lol. Thanks for the reassurance :)

OpenStudy (texaschic101):

no problem :)